Parth (parthkohli):

Solving the diophantine equation:\[a^3 + b^3 + c^3 = 100a + 10b + c \]

OpenStudy (perl):

hmmm

Parth (parthkohli):

Can you help me with that?

OpenStudy (perl):

find all integer solutions

OpenStudy (perl):

looks like we can factor something here

OpenStudy (perl):

did you try to expand (a+b+c)^3

Parth (parthkohli):

Whoo.

Parth (parthkohli):

\[a^3+3 a^2 b+3 a^2 c+3 a b^2+6 a b c+3 a c^2+b^3+3 b^2 c+3 b c^2+c^3\]

Parth (parthkohli):

So do we have to add \(3 a^2 b+3 a^2 c+3 a b^2+6 a b c+3 a c^2+3 b^2 c+3 b c^2\) to both sides?

Parth (parthkohli):

\[(a + b + c)^3 = 3 a^2 b+3 a^2 c+3 a b^2+6 a b c+3 a c^2+3 b^2 c+3 b c^2 + 100a + 10b + c \]

OpenStudy (perl):

oh look at that

OpenStudy (perl):

one second

OpenStudy (perl):

(a+b+c)^3 = a^3+3 a^2 b+3 a^2 c+3 a b^2+6 a b c+3 a c^2+b^3+3 b^2 c+3 b c^2+c^3

Parth (parthkohli):

How would you solve it now?

OpenStudy (perl):

not sure

OpenStudy (perl):

where did you get this question, is it solvable?

OpenStudy (perl):

well the simple approach is , equate terms

Parth (parthkohli):

Yes, a solution is \((1,5,3)\)

OpenStudy (perl):